What percent of 1800 is 1150? 40 is 82% of what Number?
Answer by Alan3354(69443)

What percent of 1800 is 1150?
1150/1800 = 0.63888...
=~ 63.9%
--------------
40 is 82% of what Number?
n*0.82 = 40
n = 40/0.82
n =~ 48.78